OpenStudy (anonymous):

1/i i know i is sq root of -1 but i need sm helpp

OpenStudy (anonymous):

it would be the same thing as i^-1. so if your furmiliar with the circle of i's then you would have to go backwards from i^0 which would be the same as i^3 which is equal to -i. if this doesnt make sense i can draw a picture.

OpenStudy (anonymous):

can u please :) sorry openstudy was down

OpenStudy (anonymous):

\[\frac{1}{i}=\frac{1}{i}\times \frac{i}{i}=\frac{i}{-1}=-i\]

OpenStudy (anonymous):

i have a question how come its i divided by i can u explain that to me plz